A local fire hall holds a chicken dinner and charges $8.50 for an adults dinner and $4.00 for a child's dinner. A total of 230 dinners were sold which brought in a total of $1698.50. how many adult and how many children's dinner were sold?
Let x=

Answer by ReadingBoosters(3246)

x = adults
y = child's
x+y=230
8.50x + 4y = 1698.50
8.50(230-y) + 4y = 1698.50
1955 - 8.5y + 4y = 1698.50
-4.5y = -256.5
y = 57 children's dinners sold
x = 230 - 57 = 173 adult's dinners sold
Proof
8.5(173) + 4(57) = 1470.50 + 228 = 1698.5
..................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.